Dry Cleaning Enterprises (DCE) is expected to generate $120,000 per year for the next two years. The firm will be liquidated at the end of Year 2 but no cash proceeds will be generated from the liquidation. The firm’s investors require 15% return. Management is considering what to do with the cash that will be generated at the end of the first year. It can pay the entire $120,000 in dividends (100% dividend payout policy). Alternatively, it can pay 30% in dividends at the end of Year 1 (30% payout policy) and reinvest the remaining amount in a diamond exploration project (DXP). The project is expected to generate $15,000 in cash at the end of Year 2 and these cash flows will be paid to shareholders as additional dividends. In addition, the initial investment in DXP will be recovered from liquidation at the end of Year 2.

a) What will be the firm’s value under the 100% dividend payout?

b) What will be the firm’s value under the 30% dividend policy?

c) Should DCE continue with the 100% dividend policy or change its policy to 30% dividend policy? Why?

d) DCE implements the 30% dividend policy. DCE has 3000 shares outstanding. John is a retired teacher and owns 20% of DCE. What would be John’s dividends under the 100% dividend policy?

e) Under the new dividend policy, how much cash dividends can John expect to receive at the end of Year 1 and at the end of Year 2?

f) Assume that DCE changes to 30% dividend policy. At the end of Year 1, John receives his share of dividends as you calculate in Part e. However, being retired, John would like to receive from DCE at least $24,000 at the end of Year 1. How many shares of DCE must John sell at the end of Year 1 to make his cash flow (dividends plus cash from selling shares) equal to $24,000?. Hint, John will sell the shares after he receives the dividends (on the ex-dividend day)
Answer #1

A.

Cash flows from the firm under 100% dividend policy = 120,000 per period over 3 periods

Discount rate = 15%, Number of years = 3, Payment = 120,000, Future Value = 0, Present Value = 273,987.01....

B.

30% dividend policy:

First year cash flows to shareholders = 120,000 x 0.30 = 36,000

Second year cash flow to Shareholders = 120,000 + 15,000 = 135,000

Third year cash flow to Shareholders = 120,000 + 15,000 + 84,000 = 219,000

Discount rate= 15%, Present Value = 36,000/(1.15) + 135,000/(1.15)2+ 219,000/(1.15)3= 277,379.80.....

C.

Company should change to 30% dividend policy as the change in policy would increase the value to shareholders.

D and E

First year = 36,000 x 0.20 = 7,200

Second year = 135,000 x 0.20 = 27,000

Third year = 219,000 x0.20 = 43,800.......

F.

Under the 100% dividend policy, John would have received 120,000 x 0.20 = 24,000

Under the 30% dividend policy, John would be short 24,000 - 7,200 =16,800

Share price after payment of the first dividend:

Value of the firm = 135,000/(1.15) + 219,000/(1.15)2= 282,986.77

Value per share = 282,986.77/3000 = 94.33

John must sell: 16,800 / 94.33 = 178.01 units.

i.e. 178 shares/units